For the quadratic y=(x+1) -36 or y=x+2x-35
a) give the coordinates of the intercepts of the x-axis intercept and the y-axis intercept
b) give the equation of the axis of symmetry and the co-ordinates of the turning point
please help i need guidance and an explanation looked everywhere for help most likely searching in the wrong places.
thankyou for your time tutor..

Answer by Alan3354(69443)

For the quadratic y=(x+1) -36 or y=x+2x-35
-------------
Those are not quadratic.